OpenStudy (anonymous):

u=[root2(1-i)/2]30 Find the number u.

OpenStudy (anonymous):

\[\sqrt{\frac{1-i}{2}}\]?

OpenStudy (anonymous):

where does the "30" go?

OpenStudy (anonymous):

wait,I will use equation

myininaya (myininaya):

u=\[\sqrt{\frac{(1-i)}{2}}*30\]

OpenStudy (anonymous):

\[[(\sqrt{2}(1-i)/2 ]^{30}\]

OpenStudy (anonymous):

aaah much easier!

OpenStudy (anonymous):

\[(\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i)^{30}\]

myininaya (myininaya):

\[2^\frac{30}{2}*\frac{(1-i)^{30}}{2^{30}}=2^{15}*\frac{(1-i)^{30}}{2^{30}}=\frac{(1-i)^{30}}{2^{15}}\]

OpenStudy (anonymous):

makes the idea even easier than what was written

OpenStudy (anonymous):

oh dear no my myininaya

OpenStudy (anonymous):

i bet it is what i wrote, because that is an easy one. rewrite in trig (polar, whatever) form. then multiply the angle by 30. the modulus (absolute value) is 1 so nothing to worry about since you are on the unit circle.

myininaya (myininaya):

leave me alone i don't know math

myininaya (myininaya):

its all been guesses

OpenStudy (anonymous):

wrong answer,I need to found 1-i. :S

OpenStudy (anonymous):

ok do you know how to write this in polar form?

OpenStudy (anonymous):

because once you do it is really easy

OpenStudy (anonymous):

the number is \[\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i\]

OpenStudy (anonymous):

which sits right on the unit circle

OpenStudy (anonymous):

do you prefer degrees or radians for these?

OpenStudy (anonymous):

The answer is 1-i. But I don't know how to find it

OpenStudy (anonymous):

actually that is not the answer, but we will find it. it is not hard

OpenStudy (anonymous):

but before we begin, do you know how to write this in trigonometric form?

OpenStudy (anonymous):

Do you mean A(root2/2,-root2/2) ?

OpenStudy (anonymous):

because the idea is just to multiply the angle by 30 and see what you get. the absolute value of this number is 1, and 1 to the power of 30 is 1

OpenStudy (anonymous):

no i mean as \[r(\cos(\theta)+i\sin(\theta))\]

OpenStudy (anonymous):

in this case \[r=1\] because we are right on the unit circle

OpenStudy (anonymous):

\[|\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i|=1\]

OpenStudy (anonymous):

I haven't learned anything like this yet.

OpenStudy (anonymous):

so all we need it \[\theta\] which should be easy because we see these numbers all the time

OpenStudy (anonymous):

well if you have not, you have no way to raise this to the power of 30!

OpenStudy (anonymous):

i mean you are not expected to multiply this number by itself 30 times !

OpenStudy (anonymous):

Wait,see what is extcally writes :

OpenStudy (anonymous):

ok. but if you can write in trig from raising to the power of 30 is easy. maybe this is something else

OpenStudy (anonymous):

u=(\[(\sqrt{2/2}\times w)^{30}\]

OpenStudy (anonymous):

and we know that w=1-i

OpenStudy (anonymous):

yeah well that is what i wrote

OpenStudy (anonymous):

The answer is i **

OpenStudy (anonymous):

\[(\frac{\sqrt{2}}{2}(1-i))^{30}\] like this yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well we can do it easily, but if you have not seen trig from i don't know how to explain it

OpenStudy (anonymous):

write it

OpenStudy (anonymous):

i mean you certainly are not going to multiply this number by itself 30 times

OpenStudy (anonymous):

ok this number is \[\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i\] \[=\cos(\frac{7\pi}{4})+i\sin(\frac{7\pi}{4})\]

OpenStudy (anonymous):

I have an idea. We can do this: (\frac{\sqrt{2}}{2}(1-i))^{29}*\frac{\sqrt{2}}{2}

OpenStudy (anonymous):

so when you raise to the power of 30 you just multiply the angle by 30

OpenStudy (anonymous):

I have an idea. We can do this: (\frac{\sqrt{2}}{2}(1-i))^{29}*(\frac{\sqrt{2}}{2}(1-i)

OpenStudy (anonymous):

omg!

OpenStudy (anonymous):

\[(\frac{\sqrt{2}}{2}(1-i))^{29}*\frac{\sqrt{2}}{2}(1-i)\]

OpenStudy (anonymous):

ok that is what you wrote yes?

OpenStudy (anonymous):

YE!

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

*-(-\[\sqrt{2}/2+\sqrt{2}/2\]

OpenStudy (anonymous):

ok you have to write \[\frac{7\pi}{4}\times 30=\frac{105\pi}{2}\] and realize that this is the same place on the unit circle as \[\frac{\pi}{2}\] and so the answer is \[i\]

OpenStudy (anonymous):

without polar form or trig form i really don't know how you are expected to do this problem.

OpenStudy (anonymous):

sorry that was not much help.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

myininaya (myininaya):

omg you guys made this totally harder than is should be \[(1-i)*(1-i)=1-2i-i^2=1-2i-(-1)=1-2i+1=-2i\] \[(1-i)^2=-2i\] \[(1-i)^{2*15}=(-2i)^{15}\] \[(1-i)^{30}=(-2i)^{15}\] \[(1-i)^{30}=(-2)^{15}i^{15}=-2^{15}i^{4*3+2}=-2^{15}i^2=-2^{15}(-1)=2^{15}\] \[(1-i)^{30}=2^{15}\] \[(\frac{\sqrt{2}}{2}(1-i))^{30}=\frac{2^{15}}{2^{30}}(1-i)^{30}=\frac{1}{2^{15}}*2^{15}=1\]

myininaya (myininaya):

i made a typo in first line \[(1-i)(1-i)=1-2i+i^2=1-2i+(-1)=-2i\]

myininaya (myininaya):

:)

OpenStudy (anonymous):

actually i am not kicking myself yet

myininaya (myininaya):

lol

OpenStudy (anonymous):

since this is wrong

myininaya (myininaya):

what?

OpenStudy (anonymous):

oh wait. i was looking at the first line.

OpenStudy (anonymous):

but you are right it must be a "typo"

OpenStudy (anonymous):

you do get \[(1-i)(1-i)=-2i\]

myininaya (myininaya):

yep

OpenStudy (anonymous):

the answer is i.

OpenStudy (anonymous):

everything looks good but your answer!

OpenStudy (anonymous):

because i happen to know that it is i, not 1

myininaya (myininaya):

i will look at it again

myininaya (myininaya):

but the approach looks good

OpenStudy (anonymous):

taht is ok you just forgot the \[i^{15}\] in your answer

OpenStudy (anonymous):

no it is right, you just forgot one part i think at the end. much much easier in polar form really

myininaya (myininaya):

\[i^{4*3+3}\] omg i put a 2 where that 3 was

OpenStudy (anonymous):

to hell with IE9!

myininaya (myininaya):

yes we get i :)

myininaya (myininaya):

did you see my mistake above \[(-2)^{15}*i^{15}=-2^{15}*i^{4*3+3}=-2^{15}*i^3=-2^{15}*(-i)=2^{15}*i\] \[\frac{1}{2^{15}}*(1-i)^{30}=\frac{1}{2^{15}}*(2^{15}*i)=i\]